# On Future Power Systems Digital Twins: Towards a Standard Architecture

### v1

@@ Abstract
The energy sector’s digital transformation brings mutually dependent communication and energy infrastructure, tightening the relationship between the physical and the digital world. Digital twins (DT) are the key concept for this. This paper initially discusses the evolution of the DT concept across various engineering applications before narrowing its focus to the power systems domain. By reviewing different definitions and applications, we present a new definition of DTs specifically tailored to power systems. Based on the proposed definition and extensive deliberations and consultations with distribution system operators, energy traders, and municipalities, we introduce a standard DT ecosystem architecture that offers services beyond real-time updates and can seamlessly integrate with existing transmission and distribution system operators’ processes, while reconciling with concepts such as microgrids and local energy communities based on a system-of-systems view. We also discuss the integration of power system DTs into various phases of the system’s life cycle, such as long-term planning, emphasizing challenges that remain to be addressed, such as managing measurement and model errors, and uncertainty propagation. Finally, we present our vision of how artificial intelligence (AI) and machine learning (ML) can enhance several power system DT modules established in the proposed architecture.
@@ I Introduction
The introduction of the supervisory control and data acquisition (SCADA) system and the wide area monitoring system (WAMS) are perfect examples of how digitalization in power systems has led to an increase in system reliability over the past decades [1]. Enabled by the digital paradigm shift, investments in information and communication technology (ICT) have enabled power system operators to install advanced metering infrastructure (AMI), perform real-time (partial) network monitoring, and enhance simulation accuracy [2]. Consequently, facilitating more precise predictions of the system response and thus enhancing operational efficiency [3]. However, the increasing size and complexity of power systems, the surge in distributed energy resources (DERs), and the transformation toward flexible and controllable loads require the management of immense data pools, advanced simulation and modeling efforts, and the development of more complex operation and planning approaches. A digital twin (DT) can provide a holistic approach for data processing, modeling, simulation, and service validation [4]. Thereby playing an essential role in bridging the gap between physical and digital models [5].
@@
Since then, the DT concept has been adapted to the power systems domain to evolve into a twin-centric digital control center architecture [12, 4], in which the physical system and the dynamic simulation-based digital world are interlinked by a real-time automated data flow. The capabilities of such a system include anomaly detection and threat mitigation, self-adaptation, and prediction of the system response, similar to the capabilities first envisioned in [13]. Despite years, the general DT concept in the power systems domain remains vague and surrounded by misconceptions, as exemplified in Fig. 1 [11]. For instance, digital models and shadows are often portrayed as DTs. Nevertheless, they lack one of the fundamental features of the DT concept, namely, a bidirectional real-time automated data exchange between a physical entity (or system) and its digital counterpart. In this context, the need for a customized power system DT definition has already been recognized in literature [14].
@@
This paper begins by examining the evolution of the DT concept across various engineering domains before narrowing its focus to the power systems domain. Then it identifies and highlights key features that we consider fundamental for the future development of DTs in power systems. By reviewing different definitions and capabilities, we present in Sec. II a new definition of DTs specifically tailored to power systems. Building on the proposed definition, Sec. III presents a standard DT ecosystem architecture designed to offer services beyond real-time monitoring, control, and operation. The proposed ecosystem architecture can seamlessly integrate with existing transmission system operator (TSO) and distribution system operator (DSO) processes while reconciling with concepts such as microgrids (MGs) and local energy communities (LECs) based on a system-of-systems view. This ecosystem architecture resulted from extensive deliberations and consultations with DSOs, energy traders, and municipalities, and aims to address their primary inquiries regarding the use of DTs to support their wide range of primary objectives. In Sec. IV, we elaborate on our vision for DTs in advanced power system life cycle phases (i.e., long-term planning), highlighting challenges that remain, such as managing measurement and model errors, and uncertainty propagation. Finally, in Sec. V, we present our vision of how artificial intelligence (AI) and machine learning (ML) can enhance several power system DT components and modules established in the proposed standard ecosystem architecture.
@@ II Digital Twins & Power Systems
The prospective capabilities of DTs depicted in the NASA papers [9, 10] laid the foundation for extensive research on the application of DTs and their widespread adoption in different industries. Variants of NASA’s definition have been proposed for both generic purposes and specific domains, especially in applications for manufacturing and mechatronics. The International Academy for Production Engineering (CIRP) proposes a DT definition for generic purposes that encompasses similar features to NASA’s definition. However, these features can be tailored for specific purposes [15]. By delineating a DT as a system with a wide range of tailorable features, it allows for systems ranging from a static human-triggered simulation tool to a dynamic fully autonomous forecasting tool to be labeled as a DT. Therefore, the tailorable DT definition provided by CIRP comes at the expense of the fundamental paradigm of twinning, that is, a bidirectional real-time automated data exchange between a digital and a physical entity [16].
@@ II-A Digital Twin Definitions & Features
In general, NASA’s and CIRP’s definitions highlight three key DT features: a physical entity (or system), a digital entity (or system), and their bidirectional real-time automated data exchange (see Fig. 2). Recognizing these as the main features enables DTs to offer their well-known real-time monitoring services. Nevertheless, driven by the increased complexity of engineering systems and the involved ICT infrastructure required to handle large amounts of data, additional features have been gaining more attention (see Table I). These features include services and modules that go beyond real-time monitoring, encompassing tasks such as scenario simulation, system optimization, diagnostics, and performance predictions [17, 9, 10]. Indeed, the deployment of such advanced services and modules requires standardized ways of sharing data using data interfaces as well as more complex data storage and data visualization capabilities. As a result, these data capabilities are now recognized as important DT features [18].
Despite the comprehensive list of features now recognized as fundamental for a DT, DTs are domain-specific and can have different levels of sophistication. They are generally designed to be embedded into domain-specific processes within a broader system’s life cycle (i.e., conceptualization and design, operation, planning) [19, 20, 21]. From an engineering system’s perspective, DTs are expected to provide services and support beyond real-time updates and monitoring. For instance, they can be implemented in the design phase and guide planning decisions in the long-term planning phase [22]. Due to the range of requirements in the different engineering system’s life cycle phases, we argue that the inclusion of (long-term) planning services is crucial in the DT definition (see Fig. 2). This is supported by authors in [23], that state that incorporating DTs at advanced life cycle phases (i.e., long-term decision-making) guides the design of the DT itself. Moreover, a mature DT definition facilitates the integration of such an ecosystem into already established operation and planning frameworks of any engineering system. Nevertheless, industrial applications of DTs at advanced life cycle phases remain limited or are in their conceptual phase.
@@ II-B Power System Digital Twins
In the power systems domain, DT architectures highlight similar features described in the general DT definition, i.e., a physical system, a digital system, and their interconnection [24] (see also Table I). At the TSO level, such DT features are straightforward to identify, with SCADA systems, energy management systems (EMSs), and WAMSs supporting the bidirectional data connection of assets and resources with their digital counterpart [12, 4], enabling real-time monitoring as a base for several modules. At the DSO level, the massive deployment of internet of things (IoT) sensors and smart meters facilitates data exchange, enabling real-time response for large networks [25] and advanced DER management deployment [26]. At the component level, [27, 28] developed a DT for medium voltage (MV) to low voltage (LV) transformers and dc-dc power converters, respectively. By combining high-fidelity simulation models of transformer and converter components with measurement devices, the number of measurement devices needed to support accurate real-time monitoring was minimized. The DT presented in [29] analyzes operational data in order to increase the transformer’s lifespan by ensuring safety under future operating conditions. To improve anomaly detection and maintenance scheduling, [30] develops a DT for a power plant, implementing an architecture comprised of a continuous data stream from sensors and physics-based dynamic models. In a similar power plant application, [31] develops a DT for real-time monitoring and control. Unlike [30], [31] highlights the importance of data visualization to facilitate monitoring and control by the power plant operator, although only limited visualization capabilities (i.e., industrial charts) are considered. Exploiting a more software-based architecture, [32] proposed using DTs as cloud battery management systems (BMSs) to replace the onboard BMS. This improves the computational capabilities of these systems and enables big data storage, advanced visualization, and reliable system prediction and optimization. Contrary to the perspective presented in [32], even though data storage, visualization, and interfaces are deemed necessary, they are not universally acknowledged as inherent to the DT concept, which underpins the development of the aforementioned features. Moreover, applications of DTs in long-term system planning are scarce [33].
@@ II-C A New Power System Digital Twin Definition
III Standard Digital Twin Ecosystem Architecture
We conceive the implementation of power system DTs as a comprehensive software ecosystem development and deployment challenge, necessitating a holistic approach. As a result, a standardized ecosystem architecture is needed that acknowledges the DT’s main features (see Fig. 2) and enables seamless integration into one software ecosystem. Shown in Fig. 3, and aligned with the operators’ perspective, our proposed DT ecosystem architecture is composed of five main components: (1) High-Fidelity Simulation Models, (2) Operation & Planning, (3) Grid as a Service, (4) Data Engineering, and (5) Data Analytics. The former components aim for an autonomous power system operation, enabling operators to meet their objectives from everyday system operation to long-term planning, as well as supporting external businesses and stakeholders; while the latter components are the backbone, allowing bidirectional real-time automated data exchange between the digital and physical world, as well as between the different components, modules, and models via proper data interfaces. The proposed ecosystem architecture interfaces (with appropriate data interfaces) with the TSOs’ and DSOs’ control centers. In this view, we do not propose that such a power system DT will replace the operators’ centers. Nevertheless, we envision it to evolve to be a crucial support tool.
@@ III-A High-Fidelity Simulation Models
Multiphysics, multiscale, probabilistic simulation models compose the heart of DTs and support all the capabilities of the different DT modules. From the power systems’ perspective, the High-Fidelity Simulation Models component comprises simulation models such as Power Flow Models for steady-state analysis, EMT Models for electromagnetic transients (EMT), Thermal Models for heat analysis, and Dynamic Phasor Models for analysis based on phasor measurement units (PMUs) [36]. These models are used to increase the system operators’ observability and control capabilities of their networks. Advanced simulation models can also be incorporated, such as Hardware-in-the-Loop (HiL) Models using real-time digital simulators. The literature on power system modeling and simulation tools is vast [37], and detailed taxonomies per model application are already available [38]. Due to the non-stationary environment in which a power system operates, a continuous update process is expected to be in place. This can be implemented using the Simulation Models Management module, in charge of keeping track of infrastructure (e.g., new assets) or operation changes (e.g, network topology change). Proper model version documentation must also be deployed with the support of the Data Engineering component, detailed in Sec. III-C.
@@ III-B Bidirectional Real-Time Automated Data Exchange
Digital infrastructure serves as the foundation for the digital-to-physical connection. To be fed into the DT, data is collected by field measurements, IoT devices, and smart meters from several power system assets (e.g., lines, switches, and transformers), large generation plants (e.g., nuclear plants and hydroelectric plans), customers, and DERs (e.g., solar panels, batteries, and electric vehicles). Furthermore, relevant information (e.g., meteorological data, energy market prices, and states of adjacent interconnected systems) can compose a set of contextual data, also collected and stored. In Fig. 3, although a direct data connection and exchange is depicted by the DT linking to the power system assets, such data can in practice be transferred via the proper data interface from the TSOs’ and DSOs’ WAMSs and SCADA systems, located at the operators’ premises. Besides measured data, information in the form of control commands and maintenance recommendations can also flow from the DT toward power system assets, generation plants, customers, and DERs. Nevertheless, handling large volumes of data, including structured, unstructured, and semi-structured data from various sources with different resolutions, formats, and qualities, poses a challenge [33]. Proper data management is the core function of the Data Engineering and Data Analytics components.

### v2

# On Future Power System Digital Twins: A Vision Towards a Standard Architecture

@@ Abstract
The energy sector’s digital transformation brings mutually dependent communication and energy infrastructure, tightening the relationship between the physical and the digital world. Digital twins (DT) are the key concept for this. This paper initially discusses the evolution of the DT concept across various engineering applications before narrowing its focus to the power systems domain. By reviewing different definitions and applications, we present a new definition of DTs specifically tailored to power systems. Based on the proposed definition and extensive deliberations and consultations with distribution system operators, energy traders, and municipalities, we introduce a vision of a standard DT ecosystem architecture that offers services beyond real-time updates and can seamlessly integrate with existing transmission and distribution system operators’ processes, while reconciling with concepts such as microgrids and local energy communities based on a system-of-systems view. We also discuss our vision related to the integration of power system DTs into various phases of the system’s life cycle, such as long-term planning, emphasizing challenges that remain to be addressed, such as managing measurement and model errors, and uncertainty propagation. Finally, we present our vision of how artificial intelligence and machine learning can enhance several power systems DT modules established in the proposed architecture.
@@ I Introduction
The introduction of the supervisory control and data acquisition (SCADA) system and the wide area monitoring system (WAMS) are perfect examples of how digitalization in power systems has led to an increase in system reliability over the past decades [1]. Enabled by the digital paradigm shift, investments in information and communication technology (ICT) have enabled power system operators to install advanced metering infrastructure (AMI), perform real-time (partial) network monitoring, and enhance simulation accuracy [2]. Consequently, facilitating more precise predictions of the system response and thus enhancing operational efficiency [3]. However, the increasing size and complexity of power systems, the surge in distributed energy resources (DERs), and the transformation toward flexible and controllable loads require the management of immense data pools, advanced simulation and modeling efforts, and the development of more complex operation and planning approaches. A digital twin (DT) can provide a holistic approach to data processing, modeling, simulation, and service validation [4]. Thereby playing an essential role in bridging the gap between physical and digital models [5].
@@
Since then, the DT concept has been adapted to the power systems domain to evolve into a twin-centric digital control center architecture [12, 4], in which the physical system and the dynamic simulation-based digital world are interlinked by a real-time automated data flow. The capabilities of such a system include anomaly detection and threat mitigation, self-adaptation, and prediction of the system response, similar to the capabilities first envisioned in [13]. Despite years, the general DT concept in the power systems domain remains vague and surrounded by misconceptions, as exemplified in Fig. 1 [11]. For instance, digital models and shadows are often portrayed as DTs. A simplified schematic representation of both a digital model and a digital shadow are presented in Fig. 1a and Fig. 1b, respectively. These representations lack one of the fundamental features of the DT concept, namely, a bidirectional real-time automated data exchange between a physical entity (or system) and its digital counterpart. In a digital model, once created, modifications to the physical system do not affect the digital model. Similarly, in a digital shadow, changes in the physical system influence the digital shadow, but not vice versa [11]. Moreover, the key features of DTs are domain-specific and can have different levels of sophistication. In this context, the need for a customized power system DT definition has already been recognized in literature [14]. Furthermore, researchers acknowledge the importance of DTs without achieving consensus on a universal approach, underlining the development of a standard ecosystem architecture to advance DT research and implementation [15].
@@
This paper begins by examining the evolution of the DT concept across various engineering domains before narrowing its focus to the power systems domain. Then, it identifies and highlights key features that we consider fundamental for the future development of DTs in power systems. By reviewing different definitions and capabilities, we propose a new standard definition of DTs specifically tailored to power systems. Building on the proposed definition, a vision of a standard DT ecosystem architecture designed to offer services beyond real-time monitoring, control, and operation is proposed. Later, we elaborate on our vision for DTs in advanced power system life cycle phases (i.e., long-term planning), highlighting challenges that remain, such as managing measurement and model errors, and uncertainty propagation. We also present our vision of how artificial intelligence (AI) and machine learning (ML) can enhance several power systems DT components and modules established in the proposed standard ecosystem architecture. Finally, we outline a structured approach for implementing the proposed DT ecosystem architecture. In summary, the main contributions of this paper are as follows:
  • Tailored power systems DT definition: A new standard definition of DTs in power system is proposed. The proposed DT definition extends beyond NASA’s DT definition, represented in Fig. 1c, including a long-term planning component, among others, as depicted in Fig. 1d. It underscores the importance of integrating advanced data management and high-fidelity models for power system operation and planning. Thereby aiming to consolidate DT definitions across domains to establish a standard definition for future DT implementation.
  • Standard power systems DT architecture: A standard DT ecosystem architecture is proposed to collect the author’s vision on such developments aiming to offer services beyond real-time monitoring, control, and operation. The proposed ecosystem architecture can seamlessly integrate with existing transmission system operator (TSO) and distribution system operator (DSO) processes while reconciling with concepts such as microgrids (MGs) and local energy communities (LECs) based on a system-of-systems view. This architecture resulted from extensive deliberations and consultations with DSOs, energy traders, and municipalities and aims to address their primary inquiries regarding the use of DTs to support their wide range of primary objectives. While empirical validations through case studies are crucial, the current paper serves to set forth a conceptual foundation and roadmap for the development and implementation of the proposed architecture.
  • Addressing power systems DTs as a large-scale software development challenge: Our proposed envisioned architecture addresses the large-scale software development challenges often overlooked by existing DT platforms, advocating for a standardized approach to manage the complexity and scale of modern power systems. The proposed power system DT architecture integrates advanced data management capabilities to address measurement and model errors and uncertainties, enhancing the accuracy and reliability of the DT for robust decision-making support.
  • Power systems DTs as a System-of-Systems approach: We adopt a system-of-systems perspective, allowing for integrating DTs developed at various levels, ensuring scalability and flexibility in more complex systems.
  • Future-oriented DT vision: Our proposed DT architecture incorporates AI and ML to enhance data processing, predictive maintenance, operational optimization, etc., offering advanced predictive capabilities and operational efficiencies as envisioned enhanced features of future power systems DTs platforms.
@@ II Digital Twins & Power Systems
The prospective capabilities of DTs depicted in the NASA papers [9, 10] laid the foundation for extensive research on the application of DTs and their widespread adoption in different industries. Variants of NASA’s definition have been proposed for both generic purposes and specific domains, especially in applications for manufacturing and mechatronics. The International Academy for Production Engineering (CIRP) proposes a DT definition for generic purposes that encompasses similar features to NASA’s definition. However, these features can be tailored for specific purposes [16]. By delineating a DT as a system with a wide range of tailorable features, it allows for systems ranging from a static human-triggered simulation tool to a dynamic fully autonomous forecasting tool to be labeled as a DT. Therefore, the tailorable DT definition provided by CIRP comes at the expense of the fundamental paradigm of twinning, i.e., a bidirectional real-time automated data exchange between a digital and a physical entity [17].
@@ II-A Digital Twin Definitions & Features
In general, NASA’s and CIRP’s definitions highlight three key DT features: a physical entity (or system), a digital entity (or system), and their data exchange, as can be seen in Fig. 2. Although the CIRP definition recognizes the importance of a physical-to-digital connection, that connection need not necessarily be bidirectional, real-time, and automated. Recognizing these as the main features enables DTs to offer their well-known real-time monitoring services. Nevertheless, driven by the increased complexity of engineering systems and the involved ICT infrastructure required to handle large amounts of data, additional features have been gaining more attention. The additional features outside the power systems domain are summarized in the first half of Table I and elaborated on in Sec. II-A. They include services and modules that go beyond real-time monitoring, encompassing tasks such as scenario simulation, system optimization, diagnostics, and performance predictions [18, 9, 10]. Indeed, the deployment of these advanced services and modules requires standardized ways of sharing data using data interfaces as well as more complex data storage and data visualization capabilities. As a result, these data capabilities are now recognized as important DT features [19], as shown in Fig. 2.
Despite the comprehensive list of features now recognized as fundamental for a DT, DTs are domain-specific and can have different levels of sophistication. They are generally designed to be embedded into domain-specific processes within a broader system’s life cycle (i.e., conceptualization and design, operation, planning) [20, 21, 22]. From an engineering system’s perspective, DTs are expected to provide services and support beyond real-time updates and monitoring. For instance, they can be implemented in the design phase and guide planning decisions in the long-term planning phase [23]. Due to the range of requirements in the different engineering system’s life cycle phases, we argue that the inclusion of (long-term) planning services is crucial in the DT definition and as a such, it must be included in Fig. 2). This is supported by authors in [24], that state that incorporating DTs at advanced life cycle phases (i.e., long-term decision-making) guides the design of the DT itself. Moreover, a mature DT definition facilitates the integration of such an ecosystem into already established operation and planning frameworks of any engineering system. Nevertheless, industrial applications of DTs at advanced life cycle phases remain limited or are in their conceptual phase.
@@ II-B Power System Digital Twins
In the power systems domain, DT architectures highlight similar features described in the general DT definition, i.e., a physical system, a digital system, and their interconnection [25]. See for example the CIRP and NASA definitions in Fig. 2. The DT definitions and features found in this domain are summarized in the second half of Table I and elaborated on in Sec. II-B. At the TSO level, such DT features are straightforward to identify, with SCADA systems, energy management systems (EMSs), and WAMSs supporting the bidirectional data connection of assets and resources with their digital counterpart [12, 4], enabling real-time monitoring as a base for several modules. At the DSO level, the massive deployment of internet of things (IoT) sensors and smart meters facilitates data exchange, enabling real-time response for large networks [26] and advanced DER management deployment [27]. At the component level, [28, 29] developed a DT for medium voltage (MV) to low voltage (LV) transformers and dc-dc power converters, respectively. By combining high-fidelity simulation models of transformer and converter components with measurement devices, the number of measurement devices needed to support accurate real-time monitoring was minimized. The DT presented in [30] analyzes operational data in order to increase the transformer’s lifespan by ensuring safety under future operating conditions. To improve anomaly detection and maintenance scheduling, [31] develops a DT for a power plant, implementing an architecture comprised of a continuous data stream from sensors and physics-based dynamic models. In a similar power plant application, [32] develops a DT for real-time monitoring and control. Unlike [31], [32] highlights the importance of data visualization to facilitate monitoring and control by the power plant operator, although only limited visualization capabilities (i.e., industrial charts) are considered. Exploiting a more software-based architecture, [33] proposed using DTs as cloud battery management systems (BMSs) to replace the onboard BMS. This improves the computational capabilities of these systems and enables big data storage, advanced visualization, and reliable system prediction and optimization. Contrary to the perspective presented in [33], even though data storage, visualization, and interfaces are deemed necessary, they are not universally acknowledged as inherent to the DT concept, which underpins the development of the aforementioned features. Moreover, applications of DTs in long-term system planning are scarce [34].
@@ II-C A New Power System Digital Twin Definition
III Envisioned Standard Power System Digital Twin Ecosystem Architecture
We conceive the implementation of power system DTs as a comprehensive software ecosystem development and deployment challenge, necessitating a holistic approach. As a result, a standardized ecosystem architecture is needed that acknowledges the DT’s main features already discussed and presented in Fig. 2, while enabling seamless integration into one software ecosystem. The envisioned and proposed DT ecosystem architecture, shown in Fig. 3, is developed from the operators’ perspective and aligned with their operation frameworks. Nevertheless, it is general enough to be easily extended to other applications. The envisioned architecture in Fig. 3 is composed of five main components: (1) High-Fidelity Simulation Models, (2) Operation & Planning, (3) Grid as a Service, (4) Data Engineering, and (5) Data Analytics. The former components aim for an autonomous power system operation, enabling operators to meet their objectives from everyday system operation to long-term planning, as well as supporting external businesses and stakeholders; while the latter components are the backbone, allowing bidirectional real-time automated data exchange between the digital and physical world, as well as between the different components, modules, and models via proper data interfaces. The envisioned ecosystem architecture interfaces (with appropriate data interfaces) with the TSOs’ and DSOs’ control centers. In this view, we do not foresee that such a power system DT will replace the operators’ centers. Nevertheless, we envision it to evolve to be a crucial support tool.
@@ III-A High-Fidelity Simulation Models
Multiphysics, multiscale, probabilistic simulation models compose the heart of DTs and support all the capabilities of the different DT modules. From the power systems’ perspective, the High-Fidelity Simulation Models component comprises simulation models such as Power Flow Models for steady-state analysis, EMT Models for electromagnetic transients (EMT), Thermal Models for heat analysis, and Dynamic Phasor Models for analysis based on phasor measurement units (PMUs) [37]. These models are used to increase the system operators’ observability and control capabilities of their networks. Advanced simulation models can also be incorporated, such as Hardware-in-the-Loop (HiL) Models using real-time digital simulators. The literature on power system modeling and simulation tools is vast [38], and detailed taxonomies per model application are already available [39]. Due to the non-stationary environment in which a power system operates, a continuous update process is expected to be in place. This can be implemented using the Simulation Models Management module, in charge of keeping track of infrastructure (e.g., new assets) or operation changes (e.g, network topology change). Moreover, it ensures time synchronization among the different simulation models, enabling them to cooperate and simulate using data with varying time resolutions. Proper model version documentation must also be deployed with the support of the Data Engineering component, detailed in Sec. III-C.
@@ III-B Bidirectional Real-Time Automated Data Exchange
Digital infrastructure serves as the foundation for the digital-to-physical connection. To be fed into the DT, data is collected by field measurements, IoT devices, and smart meters from several power system assets (e.g., lines, switches, and transformers), large generation plants (e.g., nuclear plants and hydroelectric plans), customers, and DERs (e.g., solar panels, batteries, and electric vehicles). Furthermore, relevant information (e.g., meteorological data, energy market prices, and states of adjacent interconnected systems) composes a set of contextual data and is also collected and stored. In Fig. 3, although a direct data connection and exchange is depicted by the DT linking to the power system assets, such data can in practice be transferred via the proper data interface from the TSOs’ and DSOs’ WAMSs and SCADA systems, located at the operators’ premises. In addition to measured data, information in the form of control commands and maintenance recommendations can also flow from the DT to power system assets, generation plants, customers, and DERs. However, handling large volumes of data, including structured, unstructured, and semi-structured data from various sources with different spatio-temporal resolutions, formats, and qualities, poses a challenge [34]. Proper data management is the core function of the Data Engineering and Data Analytics components.